Q: Suppose a firm's total cost function is TC = 16 + 5Q + 4Q2 . What is the output level that minimizes average total cost?
Answer #1

TC = 16 + 5Q + 4Q2

Dividing the whole equation by Q

ATC = TC / Q = (16 + 5Q + 4Q2 ) / Q

Differntiating TC with respect to Q to get MC

MC = 5 + 8Q

Equating MC = ATC for finding qty at which ATC minimizes

5 + 8Q = (16 + 5Q + 4Q2 ) / Q

5Q + 8Q2 = 16 + 5Q + 4Q2

4Q2 = 16

Q2 = 4

Q = 2